OpenStudy (anonymous):

I am a little confused. I need help. The question(statement) is: Write the equation in slope-intercept form of the line that has a slope of -2 and a y-intercept of 2

OpenStudy (anonymous):

ok the y intercept of 2 means that when x=0, then y=2

OpenStudy (anonymous):

basically if you have y=mx+b, and x is set to zero we remain with 2 y = 0 + b we know that at this point, y must be 2: 2 = 0 + b

OpenStudy (anonymous):



OpenStudy (anonymous):

Thanks but I already had the answer.

OpenStudy (anonymous):

what do you need help with? did you sort everything out?

OpenStudy (anonymous):

Yes I did.